# Increasing Data Rate through Shaping on Wireless Channels Subject to Mobility and Delay Spread

### v2

@@ Abstract
This letter describes how to improve performance of cellular systems by combining non-equiprobable signaling with low-density parity check (LDPC) coding for an orthogonal frequency division multiplexing system. We focus on improving performance of wireless transmission with rate $3/4$ LDPC codes employing $4$-QAM at the cell edge. We double the size of the $4$-QAM constellation by introducing a second shell of signal points, and we implement non-equiprobable signaling through a shaping code which selects the high energy shell less frequently than the low energy shell. We emphasize that the shaping code is also used to transmit information. We employ rate $1/2$ LDPC coding and select the rate of the shaping code to match that of rate $3/4$ LDPC coding using $4$-QAM. We describe how to combine coding and shaping by integrating shaping into the calculation of log-likelihood ratios (LLRs) necessary for decoding LDPC codes. We present simulation results for a representative Veh-A channel showing gains of $4$ dB at a bit error rate (BER) of $10^{-3}$. Our results suggest that it is better to increase rate at the cell edge by introducing shaping, rather than by increasing the rate of the LDPC code.
@@ I Introduction
We use the method of shaping on rings to improve 5GNR performance at the cell edge. The 5GNR standard specifies rate $1/2$ and rate $3/4$ LDPC codes which use 4-QAM at the cell edge. At a bit error rate of $10^{-3}$ the gap between the performance of the rate $1/2$ code and the rate $3/4$ code is significant - approximately $4$ dB. We bridge this gap by combining the rate $1/2$ LDPC code with a shaping code. We expand the baseline $4$-QAM constellation to the $8$-QAM constellation [^1] We refer to the expanded constellation as $8$-QAM since the expanded constellation has $8$ points. The constellation can also be viewed as a $16$-QAM constellation with the points not present in Fig. 2 carrying zero probability. shown in Fig. 2 by introducing a second shell of signal points, and we employ shaping codes that select the high energy shell less frequently than the low energy shell. The rate $3/4$ LDPC code transmits $1.5$ bits per QAM symbol, and we match this rate by introducing a shaping code with rate $0.5$ bits per QAM symbol. We ask the fundamental question of whether it is better to increase rate at the cell edge in 5GNR by introducing shaping or by increasing the rate of the LDPC code. We present simulation results for a representative Veh-A channel showing that rate $1/2$ LDPC coding with shaping improves on rate $3/4$ LDPC coding by $4$ dB at a bit error rate (BER) of $10^{-3}$. Our results demonstrate that at the cell edge where transmission rates are low, there is significant value in using shaping to contribute transmission rate. This does not appear possible with the GS and PS methods currently under consideration at 3GPP.
@@ II System Model
Consider an OFDM frame with $M$ subcarriers and $N$ symbols. Let $\Delta f$ be the subcarrier spacing. The OFDM frame occupies bandwidth $B=M\Delta f$ and time $T=N/\Delta f$. Each OFDM symbols occupies time $T_{\mathsf{s}}=1/\Delta f$. Let $\mathbf{X}\in\mathbb{C}^{M\times N}$ denote the matrix of information symbols. At the transmitter, the information symbols are converted to time domain as:
@@
$\mathbf{y}_{\mathsf{F}}=(\mathbf{I}\otimes\mathbf{F}_{M})\mathbf{H}(\mathbf{I}\otimes\mathbf{F}_{M}^{\mathsf{H}})\mathbf{x}_{\mathsf{F}}+\mathbf{w}_{\mathsf{F}}=\mathbf{H}_{\mathsf{F}}\mathbf{x}_{\mathsf{F}}+\mathbf{w}_{\mathsf{F}},$ (10)
@@ II-A Channel estimation
In practice, systems are required to estimate the channel at the receiver. To estimate the channel, a reference symbol or pilot is transmitted. In this paper, we consider the Kronecker pilot pattern (also called the Type-2 demodulation reference signal (DMRS)) [1, 2]. In this setting pilots are placed on specific subcarriers in an OFDM symbol. The estimates at the pilot locations are obtained using a least squares estimator [11] and to get the estimate at the rest of the locations a two-dimensional interpolation is carried out. For the two-dimensional interpolation, the (pre-computed) time and frequency covariance matrices are used, which are computed as:
$\mathbf{C}_{\mathsf{t}}=\mathbb{E}[\mathsf{diag}(\mathbf{H}_{\mathsf{F}}^{\mathsf{H}})\mathsf{diag}(\mathbf{H}_{\mathsf{F}})],\mathbf{C}_{\mathsf{f}}=\mathbb{E}[\mathsf{diag}(\mathbf{H}_{\mathsf{F}})\mathsf{diag}(\mathbf{H}_{\mathsf{F}}^{\mathsf{H}})],$ (11)
respectively [^3] Here, we consider the diagonal values when we compute the covariance matrices. This is because under fractional delay-Doppler channel, the channel matrix is not strictly diagonal due to inter-carrier interference (ICI) in OFDM. Taking the whole matrix introduces ICI effects into the time and frequency covariance matrix.. A linear minimum mean squared error (LMMSE) interpolator is used for the two-dimensional interpolation of the channel values [6] which gives the estimate of the diagonals of $\mathbf{H}_{\mathsf{F}}$. This forms the estimated channel matrix which is used for equalization.
@@ III-A Shaping Encoder
Figure 2: Constellation points used for communication in the shaping scheme.
At the transmitter, an augmented constellation is generated. An example is shown for 4-QAM, in Fig. 2. The augmented constellation is a combination of an inner (or the original) constellation and an outer constellation [^4] In this paper, although we divide the constellation into two regions, inner and outer, the shaping idea is very general. For example, the constellation could be divided into multiple partitions as shown in [4].. Each symbol from the inner constellation is mapped to a point on the outer constellation. Such pairs have the same bit representation. In the example, if the point $1+1j$ is represented by the binary vector $0\ 0]^{\top}$ then so is the point 3+1j-3+1j. In the Figure, each constellation point on the 4-QAM shown by a square, circle, triangle, or star, is mapped to a point on the outer constellation represented by the corresponding shaded shape. The choice between the inner and outer constellation is provided by the shaping code. Note that the shaping code is not actually transmitted, the extra information bits are conveyed through power during transmission.
Figure 3: Block diagram showing a shaping encoder. Uncoded bits are provided as input. A shaping code $1\ 0\ 0]^{\top}$ (length 33) is used.
A shaping code is a collection of binary vectors of length $z$ which have a certain sparsity. For example, a $1$-sparse shaping code of length $z$ would have $z+1$ vectors, where $z$ vectors have sparsity $1$ and one vector has 0 sparsity. The constellation symbols that are to be transmitted are divided into vectors $\mathbf{s}$ of length $z$ at the transmitter. The shaping encoder picks a $z$-length codeword, say $\mathbf{c}$, at random from the codebook. Let $f$ be a function that maps each constellation point in the inner constellation to the corresponding point on the outer constellation. Then the operation performed by the shaping encoder is given by:
@@
for $i=0,1,\cdots,z-1$ and $\bar{\mathbf{s}}\in\mathbb{C}^{z\times 1}$ is the encoded $z$-length vector that is transmitted. The encoder transmits information symbols from the inner constellation whenever the shaping code at the corresponding index is 0 and from the outer constellation otherwise. Note that, the all 0 shaping codeword corresponds to no shaping.
The encoder operation is shown in Fig. 3. For the example, the $z=3$ and the shaping code is $1\ 0\ 0]^{\top}$. A point from the outer constellation is transmitted at the first location and points from the inner constellation are transmitted at the second and third locations.
@@
III-B Shaping Decoder
$\mathbf{d}=\underset{s\in\mathcal{S}}{\min}\ \|\mathbf{y}-s\|,$ (13)
The shaping codeword is decoded as:
[…]
However, since the shaping codeword has a pre-defined structure (for example, sparsity) the decoded codeword may not always be a valid codeword. To get to the valid codeword, we use the fact that while decoding constellation symbols, received symbols closer to the decision boundary are less reliable than the ones closer to the constellation points. The less reliable received symbols need to be moved a greater distance to be decoded to the nearest codeword compared to the more reliable ones. The less-reliable symbols are therefore more likely to be in error. This motivates the following decoding rule. When the codeword is not valid, the index of the “dirtiest” symbol (i.e., the symbol that needs to be moved a great distance for decoding) is computed as:
$i=\underset{i=0,1,\cdots,z-1}{\arg\max}\ \mathbf{d}[i],$ (17)
and the bit at $\hat{\mathbf{c}}[i]$ is flipped (since it is the least reliable). If this results in a valid codeword, the process is stopped else the process is continued to find the next dirtiest symbol index and so on. This procedure gives the decoded shaping codeword.
III-C LDPC Decoding
@@
Figure 4: Coded bit-error performance comparison between OFDM modulation with and without shaping. 44-QAM, $15$-length shaping with 11 sparsity.
III-D Rate, PAPR, and Complexity Considerations
@@
For example, for a 4-QAM system in Fig. 2, with $3$-length shaping with sparsity $1$, $P_{\mathsf{avg}}=0.75\cdot 2+0.25\cdot 10=4$ and the $\mathrm{PAPR}=2.5$.
At the transmitter, to encode the shaping codeword, the transmitter needs to implement (12), which is a look-up table based operation and does not incur computational complexity. At the receiver, the minimum distance (in (15)) needs to be computed for all the constellation points (twice as many times without shaping) and incurs additional complexity $\mathcal{O}(|\mathcal{S}_{\mathsf{o}}|)$. In the LLR computation, the complexity involving combining the LLRs incurs complexity $\mathcal{O}(|\mathcal{S}_{\mathsf{o}}|)$ (since insider the for loop, all computations can be carried out in one clock cycle and incur complexity $\mathcal{O}(1)$). The additional complexity introduced by shaping is only linear in the number of additional constellation points and does not scale with the frame dimensions.
Figure 5: Coded bit-error performance comparison between OFDM modulation with and without shaping. 44-QAM, $23$-length shaping with 33 sparsity.
@@ IV Numerical Results
In this Section we present the numerical results evaluating the performance of the OFDM scheme with shaping. For all the results presented here, we consider a practical setting where estimated channel is used (per Sec. II-A) for equalization along with LDPC encoding. We consider an OFDM system with $M=72,N=14,\Delta f=30$ kHz. For all the simulations we consider the Vehicular-A channel model [9] with a maximum Doppler spread $\nu_{\max}$ of $815$ Hz. Each path Doppler is obtained using the Jake’s spectrum $\nu_{p}=\nu_{\max}\cos(\theta)$, where $\theta\sim\mathsf{U}[-\pi,\pi)$. We use the 5G LDPC encoder and decoder provided by Sionna [7]. The message length $k$ and code length $n$ are specified for each code in the following subsections.
@@ V Conclusions
We have shown that non-equiprobable signaling provides superior BER performance on wireless channels subject to mobility and delay spread. The advantages are particularly striking for small constellations such as $4$-QAM, and we have presented simulations for a representative Veh-A channel showing gains of $4$ dB at a bit error rate of $10^{-3}$. We have described a method of shaping on rings that can be applied to standard QAM constellations, and is able to achieve any desired fractional bit rate. We have described how to combine LDPC codes with shaping, by integrating shaping into the calculation of log-likelihood ratios (LLRs) necessary for LDPC decoding.

### v3

@@ Abstract
This letter describes how to improve performance of cellular systems by combining non-equiprobable signaling (shaping) with low-density parity check (LDPC) coding for an orthogonal frequency division multiplexing system. We focus on improving performance at the cell edge, where the 5G standard specifies a suite of LDPC codes with different rates that are applied to 4-QAM. We employ the method of shaping on rings which adds to the transmission rate as it shapes the input distribution. We double the size of the $4$-QAM constellation by introducing a second shell of signal points, and we implement non-equiprobable signaling through a shaping code which selects the high energy shell less frequently than the low energy shell. We describe how to combine coding and shaping by integrating shaping into the calculation of log-likelihood ratios (LLRs) necessary for decoding LDPC codes. We employ rate $1/2$ LDPC coding and select the rate of the shaping code to match that of rate $3/4$ LDPC coding using $4$-QAM. We present simulation results for a representative Veh-A channel showing gains of $4$ dB at a bit error rate (BER) of $10^{-3}$. When we choose an LDPC code from the 5G suite to match the BER performance of rate $1/2$ LDPC coding with shaping we show that transmission rate can be improved by 2020%.
@@ I Introduction
We use the method of shaping on rings to improve 5GNR performance at the cell edge. The 5GNR standard specifies rate $1/2$ and rate $3/4$ LDPC codes which use 4-QAM at the cell edge. At a bit error rate of $10^{-3}$ the gap between the performance of the rate $1/2$ code and the rate $3/4$ code is significant - approximately $4$ dB. We bridge this gap by combining the rate $1/2$ LDPC code with a shaping code. We expand the baseline $4$-QAM constellation to one of the $8$-QAM constellations [^1] We refer to both constellations as 8-QAM since each has 8 signal points, and we have observed no difference in performance between the two constellations. shown in Fig. 2 by introducing a second shell of signal points, and we employ shaping codes that select the high energy shell less frequently than the low energy shell. The rate $3/4$ LDPC code transmits $1.5$ bits per QAM symbol, and we match this rate by introducing a shaping code with rate $0.5$ bits per QAM symbol. We ask the fundamental question of whether it is better to increase rate at the cell edge in 5GNR by introducing shaping or by increasing the rate of the LDPC code. We present simulation results for a representative Veh-A channel showing that rate $1/2$ LDPC coding with shaping improves on rate $3/4$ LDPC coding by $4$ dB at a bit error rate (BER) of $10^{-3}$. Our results demonstrate that at the cell edge where transmission rates are low, there is significant value in using shaping to contribute transmission rate. This does not appear possible with the GS and PS methods currently under consideration at 3GPP.
@@ II System Model
Consider an OFDM frame with $M$ subcarriers and $N$ symbols. Let $\Delta f$ be the subcarrier spacing. The OFDM frame occupies bandwidth $B=M\Delta f$ and time $T=N/\Delta f$. Each OFDM symbols occupies time $T_{\mathsf{s}}=1/\Delta f$. The system model can be represented as (see [10] for a detailed derivation of the system model): Let $\mathbf{X}\in\mathbb{C}^{M\times N}$ denote the matrix of information symbols. At the transmitter, the information symbols are converted to time domain as:
@@
$\mathbf{y}_{\mathsf{F}}=\mathbf{H}_{\mathsf{F}}\mathbf{x}_{\mathsf{F}}+\mathbf{w}_{\mathsf{F}}.$ (10)
@@ II-A Channel estimation
In practice, systems are required to estimate the channel at the receiver. To estimate the channel, a reference symbol or pilot is transmitted. In this paper, we consider the Kronecker pilot pattern (also called the Type-2 demodulation reference signal (DMRS)) [1, 2]. In this setting pilots are placed on specific subcarriers in an OFDM symbol. The estimates at the pilot locations are obtained using a least squares estimator [12] and to get the estimate at the rest of the locations a two-dimensional interpolation is carried out. For the two-dimensional interpolation, the (pre-computed) time and frequency covariance matrices are used [^3] Here, we consider the diagonal values when we compute the covariance matrices. This is because under fractional delay-Doppler channel, the channel matrix is not strictly diagonal due to inter-carrier interference (ICI) in OFDM. Taking the whole matrix introduces ICI effects into the time and frequency covariance matrix., which are computed respectively as:
$\mathbf{C}_{\mathsf{t}}=\mathbb{E}[\mathsf{diag}(\mathbf{H}_{\mathsf{F}}^{\mathsf{H}})\mathsf{diag}(\mathbf{H}_{\mathsf{F}})],\mathbf{C}_{\mathsf{f}}=\mathbb{E}[\mathsf{diag}(\mathbf{H}_{\mathsf{F}})\mathsf{diag}(\mathbf{H}_{\mathsf{F}}^{\mathsf{H}})].$ (11)
A linear minimum mean squared error (LMMSE) interpolator is used for the two-dimensional interpolation of the channel values [6] which gives the estimate of the diagonals of $\mathbf{H}_{\mathsf{F}}$. This forms the estimated channel matrix which is used for equalization.
@@ III-A Shaping Encoder
Figure 2: Two 88-QAM constellations with identical performance.
We augment the original $4$-QAM constellation to obtain the $8$-QAM constellations shown in Fig. 2. Points in the augmented constellation are labeled $\mathbf{A}$, $\mathbf{B}$, $\mathbf{C}$, or $\mathbf{D}$ with different labels corresponding to different residue classes modulo $4$. Every label appears once in the inner constellation and once in the outer constellation. The choice between the inner and outer constellation is provided by the shaping code. Note that the shaping code is not actually transmitted, the extra information bits are conveyed through power during transmission.
Figure 3: The shaping codeword [1 0 0][1\ 0\ 0] transforms a sequence [ABC][A\ B\ C] of labels to a sequence $\mathbf{A}\ B\ C]$ of transmitted signal points.
An $l$-sparse shaping code of length $z$ is a collection of binary vectors with at most $l$ entries equal to $1$. The LDPC code selects a sequence of $z$ 44-QAM points (labels). The labels are input to the shaping code which chooses between the representatives in the inner and outer constellations. If $f$ is the function that maps an inner constellation point to the corresponding outer point, then the shaping code performs the following function.
@@
for $i=0,1,\cdots,z-1$ and $\bar{\mathbf{s}}\in\mathbb{C}^{z\times 1}$ is the encoded $z$-length vector that is transmitted. The encoder transmits information symbols from the inner constellation whenever the shaping code at the corresponding index is 0 and from the outer constellation otherwise. Note that, the all 0 shaping codeword corresponds to no shaping (see Fig. 3).
@@
III-B Shaping Decoder
$\mathbf{d}=\underset{s\in\mathcal{S}}{\min}\ \|\mathbf{y}-s\|,$ (13)
The shaping codeword is decoded as:
[…]
However, since the shaping codeword has a pre-defined structure (for example, sparsity) the decoded codeword may not always be a valid codeword. If the estimate $\hat{\mathbf{c}}$ is not a shaping codeword we change one or more entries to produce a shaping codeword. If $\mathbf{d}[i]$ is the distance between the $i$th received symbol and the $i$th symbol of $\hat{\mathbf{c}}$, then larger metrics $\mathbf{d}[i]$ correspond to less reliable symbols. We identify the symbol with largest metric and flip that symbol (replacing an outer symbol with the paired inner symbol or vice versa). If this results in a valid codeword, the process is stopped else the process is continued to find the next dirtiest symbol index and so on. This procedure gives the decoded shaping codeword.
III-C LDPC Decoding
@@
III-D Rate and PAPR
@@
For the 4-QAM system in Fig. 2, with $3$-length shaping with sparsity $1$, $P_{\mathsf{avg}}=0.75\cdot 2+0.25\cdot 10=4$ and the $\mathrm{PAPR}=2.5$.
Figure 4: The rate $1/2$ LDPC code with a $1$-sparse length $15$ shaping code has the same transmission rate as the rate $5/8$ LDPC code, but improves performance by $1.5$ dB at a BER of $10^{-2}$.
@@ IV Numerical Results
In this Section we present the numerical results evaluating the performance of the OFDM scheme with shaping. For all the results presented here, we consider a practical setting where estimated channel is used for equalization along with LDPC encoding. We consider an OFDM system with $M=72,N=14,\Delta f=30$ kHz. For all the simulations we consider the Vehicular-A channel model [9] with a maximum Doppler spread $\nu_{\max}$ of $815$ Hz. Each path Doppler is obtained using the Jake’s spectrum $\nu_{p}=\nu_{\max}\cos(\theta)$, where $\theta\sim\mathsf{U}[-\pi,\pi)$. We use the 5G LDPC encoder and decoder provided by Sionna [7]. The message length $k$ and code length $n$ are specified for each code in the following subsections.
@@ IV-C $16$-QAM, $23$-length shaping
Figure 6: The rate $1/2$ LDPC code with a $3$-sparse length $23$ shaping code has the same transmission rate as the $3/4$ LDPC code, but improves performance by $4$ dB at a BER of $10^{-3}$. It has slightly better BER performance than the rate $5/8$ LDPC code but improves transmission rate by 2020%.
@@ V Conclusions
We aimed to improve performance at the cell edge, where the 55G wireless standard specifies a suite of LDPC codes with different rates that are applied to $4$-QAM. We have demonstrated that the combination of rate $1/2$ LDPC coding and shaping improves upon the performance of this baseline suite of codes. The rate $1/2$ LDPC code with a $3$-sparse length $23$ shaping code has the same transmission rate as the rate $3/4$ LDPC code, but improves performance by $4$ dB at a BER of $10^{-3}$. The same combination has slightly better BER performance than the rate $5/8$ LDPC code but improves transmission rate by 2020%. Our results suggest that at the cell edge where transmission rates are low, there is significant value in using shaping to contribute transmission rate. When decoding the combination of LDPC codes and shaping codes, we have described a method of calculating LLRs that may be of independent interest.